OpenStudy (anonymous):

A music store sold 1,112 music CDs one month. This was 20% less than the number of music CDs the store sold the previous month. The following equation represents this situation, where x represents the number of music CDs sold the previous month. x − 0.2x = 1,112 How many music CDs did the store sell in both months combined?

OpenStudy (anonymous):

I need help with this one @dmezzullo

OpenStudy (anonymous):

:)

OpenStudy (dmezzullo):

brb puppy

OpenStudy (anonymous):

k

OpenStudy (dmezzullo):

i multiplied and got an answer but i am not sure about it

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

I'll tell u the answer n then u work it out....it is 2502

OpenStudy (anonymous):

That's not the right answer though @Harkirat

OpenStudy (anonymous):

What answer hv u got as "right answer"??

OpenStudy (anonymous):

Well I didn't get the right answer, I checked your answer, and it's not right, because: 2502 - 0.2(2502) = 1,112 2502 - 501 = 1,112 2,004 = 1,112 That's not right.

OpenStudy (anonymous):

Your working is completely wrong. I gave you the total sale of CDs of the two months combined because that is the final answer for this question.....

OpenStudy (anonymous):

x − 0.2x = 1,112 0.8x = 1112 divide both sides by 0.8, and u get x = 1390 so the sales for the previous month were 1390 CDs. Total of both months is 1390+1112=2502

OpenStudy (anonymous):

wait. where did u get 0.8?

OpenStudy (anonymous):

x means 1x so 1x - 0.2 x = (1-0.2)x = 0.8x

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

I mean x = 1x. but (1-.02)x doesn't = 0.8x

OpenStudy (anonymous):

What will you get if you subtract 0.2 from 1?????

OpenStudy (anonymous):

You can see I am right... If the sales of first month are 1390 Then sales of next month are 1390 - 20% of 1390 = 1390 - 20/100 * 1390 = 1390 - 0.2*1390 = 1390 - 278 = 1112 the sales for second month

OpenStudy (anonymous):

ooooooooooooooooooooooooooooooooooooooooooooooooh oooooopssiieeee myyyy baaaaaaaaaaaaaad

OpenStudy (anonymous):

so the combined sales of the two months ARE 2502

OpenStudy (anonymous):

Yeah. :3 sorrrryyyy I was having a blonde moment, and thinking I was right. Thanks Harkirat! :)

OpenStudy (anonymous):

No worries, now that you understand it☺